40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 29 11:51:27 2021, Anchal, Summary, LSC, Start of measuring IMC WFS noise contribution in ar cavity length noise 
    Reply  Thu Apr 29 15:11:33 2021, gautam, Update, CDS, RFM  RFM_errs.pngScreenshot_2021-04-29_15-12-56.png
       Reply  Thu Apr 29 17:43:16 2021, Koji, Update, CDS, RFM  
       Reply  Thu Apr 29 17:43:48 2021, Anchal, Update, CDS, F2A Filters double check F2AFiltersON.png
          Reply  Fri Apr 30 00:20:30 2021, gautam, Update, CDS, F2A Filters double check 
    Reply  Thu Apr 29 17:51:19 2021, Anchal, Summary, LSC, Start of measuring IMC WFS noise contribution in arm cavity length noise WFS1_PIT_exc_80-100Hz_Arms_ASD.pdf
       Reply  Fri Apr 30 00:18:40 2021, gautam, Summary, LSC, Start of measuring IMC WFS noise contribution in arm cavity length noise 
       Reply  Mon May 3 09:14:01 2021, Anchal, Paco, Update, LSC, IMC WFS noise contribution in arm cavity length noise WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdfWFS1_PIT_Noise_Inj_Test_ARM_locked.pdf
          Reply  Mon May 3 17:28:58 2021, Anchal, Paco, Rana, Update, LSC, IMC WFS noise contribution in arm cavity length noise excitationoftheMCanglessothatwecanseesomethingdotpng.pngexcitationoftheMCanglessothatwecanseesomethingdotpngbutthistimeitsMC2.png
             Reply  Tue May 4 11:43:09 2021, Anchal, Paco, Update, LSC, IMC WFS noise contribution in arm cavity length noise ArmCavNoiseContributions.pdfIOO-MC1_PIT_NoiseInjTest2.pdfIOO-MC1_PIT_NoiseInjTest_AWGGUI_Config.png
                Reply  Fri May 7 11:54:02 2021, Anchal, Paco, Update, LSC, IMC WFS noise contribution in arm cavity length noise ITMX-XARM_TF.pdfITMY-YARM_TF.pdfArmCavNoiseContributions.pdf
                   Reply  Mon May 10 18:19:12 2021, Anchal, Paco, Update, LSC, IMC WFS noise contribution in arm cavity length noise, Corrections ArmCavNoiseContributions.pdf
                      Reply  Wed May 12 10:53:20 2021, Anchal, Paco, Update, LSC, PSL-IMC PDH Loop and XARM PDH Loop diagram IMC_SingleArm.pdf
Message ID: 16097     Entry time: Thu Apr 29 15:11:33 2021     In reply to: 16095     Reply to this: 16099   16100
Author: gautam 
Type: Update 
Category: CDS 
Subject: RFM  

The problem here was that the RFM errors cropped up again - seems like it started ~4am today morning judging by TRX trends. Of course without the triggering signal the arm cavity couldn't lock. I rebooted everything (since just restarting the rfm senders/receivers did not do the trick), now arm locking works fine again. It's a bit disappointing that the Rogue Master setting did not eliminate this problem completely, but oh well...

It's kind of cool that in this trend view of the TRX signal, you can see the drift of the ETMX suspension. The days are getting hot again and the temp at EX can fluctuate by >12C between day and night (so the "air-conditioning" doesn't condition that much I guess 😂 ), and I think that's what drives the drift (idk what the transfer function to the inside of the vacuum chamber is but such a large swing isn't great in any case). Not plotted here but i hypothesize TRY levels will be more constant over the day (modulo TT drift which affects both arms).

The IMC suspension team should double check their filters are on again. I am not familiar with the settings and I don't think they've been added to the SDF.

Attachment 1: RFM_errs.png  601 kB  Uploaded Thu Apr 29 16:12:32 2021  | Show | Hide all | Show all
Attachment 2: Screenshot_2021-04-29_15-12-56.png  31 kB  Uploaded Thu Apr 29 16:13:02 2021  | Hide | Hide all | Show all
Screenshot_2021-04-29_15-12-56.png
ELOG V3.1.3-